A firm has a profit margin of 4.4% and an equity multiplier of 2.3. Its sales are $143 million, and it has total assets of $95 million. What is its ROE?

Carry out calculations to at least 4 decimal places. Enter percentages as whole numbers. Example: 3.03% should be entered as 3.03. Do not include commas or dollar signs in numerical answers.
Answer #1

Total asset turnover=Sales/Total assets

=(143/95)=1.5053(Approx)

ROE=Profit margin*Total asset turnover*Equity multiplier

=(0.044*2.3*1.5053)

which is equal to

=15.2333(Approx).